the value of observed and calculated molecullar wts of silver nitrate are92.64 and 170 respectively.The degree of dissociation would be

Grade:12

1 Answers

Karam singh
5 years ago
i=calculated colligative property/observed colligative propertyi=170/92.64=1.835For AgNO3 (n)=2i=1+alpha(n-1)alpha(degree of dissociation)=i-1=1.835-1=0.835